# Moments and the Basic Equilibrium Equations

### v3

@@ 4. The space of weighted systems with $\mu_{1}={\overrightarrow{O}}$
From now on we let $X=\{x_{1},\ldots,x_{n}\}$. Denote the dimension of the affine closure of $X$ by $\text{dim}(X)$, or just $d$, and call it the dimension of $X$. The difference between the maximum possible value for $\text{dim}(X)$, which is n1n-1, and $d$, that is c=(n1)dc=(n-1)-d, is called the codimension of $X$, and we write $\text{codim}(X)=c$. A weight function $w:X\longrightarrow\mathbb{R}$ may alternatively be represented by the vector $W\in\mathbb{R}^{n}$ whose coordinates are $w(x_{j})=w_{j}$.
Hereafter we shall assume that $X$ is a subset of $\mathbb{R}^{d}$. Select a reference frame at the origin $O\in\mathbb{R}^{d}$, identify points and vectors through $p\leftrightarrow\overrightarrow{Op}$, and use coordinates to represent points and vectors alike. Consider the (augmented) configuration matrix of $X$,
$\mathcal{X}=\left[\begin{array}[]{cccc}1&1&\cdots&1\\ x_{1}&x_{2}&\cdots&x_{n}\end{array}\right].$
Regard $\mathcal{X}$ as a linear map from $\mathbb{W}=\mathbb{R}^{n}$ to $\mathbb{R}^{d+1}$, where $\mathbb{W}$ consists of all possible weight functions (in vector form) defined on $X$. We call $\mathbb{W}$ the weight space of $X$. The kernel of $\mathcal{X}$ is the subspace consisting of weight vectors such that $\mu_{0}=0$ and $\mu_{1}=\overrightarrow{O}$, that is, the subspace
$\mathbb{W}_{0}(X)=\{W\in\mathbb{W}\,\big/\,\ \mu_{1}=\overrightarrow{O}\}.$
The main objective in this section is to describe $\mathbb{W}_{0}(X)$.
@@
The configuration matrix $\mathcal{X}$ has the same rank as the matrix
@@
that is, $d+1$. The nullity of $\mathcal{X}$ is thus n(d+1)=(n1)d=cn-(d+1)=(n-1)-d=c. This is the dimension of $\mathbb{W}_{0}(X)$. ∎
@@
Recall $\bigwedge$ represents the exterior product of vector spaces. If a vector space has finite dimension, then its $k$-th exterior power is equal to the trivial space for all $k$ greater than the dimension of the vector space [20]. ∎
Many identities involving the values of distinct weight functions arise from theorem 4.2. This topic will be further discussed in section 5.
Next we present a systematic method for constructing bases for $\mathbb{W}_{0}(X)$.
As we shall soon see, a key role is played by configurations whose codimension is 1. Borrowing the terminology from Celestial Mechanics, such a configuration will be designated as Dziobek configuration, or Dziobek subconfiguration, when it is a proper subset of a larger configuration.
Notice that a configuration $X$ has dimension $d$ if and only if $X$ contains an ascending chain of simplexes $S_{0}\subset\ldots\subset S_{d-1}\subset S_{d}$, where each $S_{k}$ is $k$-dimensional, $k=0,\ldots,d-1,d$. We refer to any simplex of dimension $d$ contained in $X$ as a core of $X$. There are n(d+1)=cn-(d+1)=c points in $X\setminus S_{d}$, so we have exactly $c$ distinct Dziobek subconfigurations $D^{(1)},\ldots,D^{(c)}$ containing the core $S_{d}$. Figure 1 depicts one such arrangement of simplexes and Dziobek subconfigurations of $X$. We call each such diagram a Dziobek tree for $X$.
Consider a Dziobek tree for $X$. If $c=0$, we have that $X$ is the core $S_{d}$ and, by proposition 4.1, $\mathbb{W}_{0}(X)$ is the trivial space. The cases $c>0$ are dealt with in the next proposition. For each Dziobek subconfiguration $D^{(l)}$ containing $S_{d}$, let us embed the corresponding $\mathbb{W}_{0}(D^{(l)})$, $l=1,\ldots,c,$ as a (one-dimensional) subspace of $\mathbb{W}_{0}(X)$ through the map
(4.1) $\iota_{l}:\mathbb{W}_{0}(D^{(l)})\hookrightarrow\mathbb{W}_{0}(X),$
which is defined by placing zeros at the coordinates corresponding to the n(d+2)=c1n-(d+2)=c-1 points in $X\setminus D^{(l)}$.
Figure 1. Sketch of a Dziobek tree. The “trunk” is the ascending chain of $k$-dimensional simplexes $S_{k}$, $k=0,\ldots,d$, and the “canopy” consists of all Dziobek subconfigurations $D^{(l)}$, $l=1,\ldots,c$, containing the core $S_{d}$.
Proposition 4.3.
To every core $S_{d}$ of a configuration $X$ of codimension $c>0$ there corresponds a unique direct sum decomposition
$\mathbb{W}_{0}(X)=\bigoplus_{l=1}^{c}\iota_{l}\left(\mathbb{W}_{0}(D^{(l)})\right),$
where $D^{(l)}$, $l=1,\ldots,c$, are the Dziobek subconfigurations of $X$ which contain $S_{d}$.
Proof.
After possibly reindexing the elements of $X$, we may assume that the core is $S_{d}=\{x_{1},\ldots,x_{d+1}\}$. Let $D^{(l)}=S_{d}\cup\{x_{d+l+1}\}$, $l=1,\ldots,c$. $D^{(l)}$ has codimension 1, so $\mathbb{W}_{0}(D^{(l)})$ is spanned by a weight vector $(w_{1},\ldots,w_{d+1},w_{d+l+1})$ such that
(4.2) $\sum_{j=1}^{d+1}w_{j}+w_{d+l+1}=0,\qquad\sum_{j=1}^{d+1}w_{j}x_{j}+w_{d+l+1}x_{d+l+1}=O.$
We must have that $w_{d+l+1}\neq 0$, for all $l=1,\ldots,c$, otherwise the above equations would provide a nontrivial linear combination of the vectors $\overrightarrow{x_{1}x_{j}}$, $j=2,\ldots,d+1$, contradicting the hypothesis that $S_{d}$ is a core. By adding c1c-1 zeros as previously explained, we embed each of the spanning vectors for the Dziobek subconfigurations $D^{(l)}$ into $\mathbb{W}_{0}(X)$, thus obtaining the collection of $c$ weight vectors
$(*,w_{d+2},0,\ldots,0),\ (*,0,w_{d+3},\ldots,0),\ \ldots,\ (*,0,\ldots,0,w_{n}),$
where each * represents a (nonzero) vector in $\mathbb{R}^{d+1}$. This collection is clearly a basis for $\mathbb{W}_{0}(X)$. ∎
Remark 4.4.
Equations (4.2) characterize $w_{1},\ldots,w_{d+1}$ as the homogeneous barycentric coordinates of $x_{d+l+1}$ with respect to the simplex $S_{d}$, that is to say, $x_{d+l+1}=\text{bar}\left(S_{d},w\big|_{S_{d}}\right)$. Indeed, equations (4.2) imply that
$(w_{1}+\ldots+w_{d+1})x_{d+l+1}=w_{1}x_{1}+\ldots w_{d+1}x_{d+1},$
with $w_{1}+\ldots+w_{d+1}\neq 0$. Compare with part (2) of proposition 2.3.
Considering all possible Dziobek trees for $X$, we see from proposition 4.3 that the maximum number of bases obtainable by our method is ${n\choose d+1}={n\choose c}$. The maximum is attained when each subconfiguration of $X$ formed by $d+1$ points is a core of $X$. In this case, we call $X$ a generic configuration. A configuration is termed nongeneric if it is not generic. The type of a nongeneric configuration refers to the number of its subconfigurations with $d+1$ points that are not cores. We illustrate the possibilities for five points on a plane in figure 2. An application is made in example 5.3.
Figure 2. Planar configurations of five points: generic (left), nongeneric of type 11 (left center), nongeneric of type 2 (center right), and nongeneric of type 44 (right).
For the case $c=1$, a generator with geometric meaning for $\mathbb{W}_{0}(X)$ can be obtained using basic properties of determinants. The following proposition is proved in [28].
Proposition 4.5.
Let $X=\{x_{1},\ldots,x_{n}\}$, and let $X_{j}$ be the subconfiguration $X\setminus\{x_{j}\}$. Denote by $\mathcal{X}_{j}$ the configuration matrix of $_{j}. XX$ has codimension $1$ if and only if $\mathbb{W}_{0}(X)$ is spanned by the weight vector
$W=(-\Delta_{1},\Delta_{2},-\Delta_{3},\ldots,(-1)^{n}\Delta_{n}),$
where $\Delta_{j}=\det(\mathcal{X}_{j})$.
The cases $c>1$ require a definition and additional notation.
The (oriented) volume of a configuration $X$ of codimension $c$ is the determinant $\det(\mathcal{X})$ of its configuration matrix $\mathcal{X}$ if $c=0$, and zero otherwise. Let $X_{j_{1}\ldots j_{\beta}}$ denote the subconfiguration $X\setminus\{x_{j_{1}},\ldots,x_{j_{\beta}}\}$, where $j_{1}<\ldots<j_{\beta}$, and let $\mathcal{X}_{j_{1}\ldots j_{\beta}}$ be the configuration matrix of $X_{j_{1}\ldots j_{\beta}}$. Denote the volume of $X_{j_{1}\ldots j_{\beta}}$ by $\Delta_{j_{1}\ldots j_{\beta}}$. The construction of bases for $\mathbb{W}_{0}(X)$ using the volume barycentric coordinates $\Delta_{j_{1}\ldots j_{\beta}}$ is similar to the one done in the proof of proposition 4.3 and employs the embeddings $\iota_{l}$, see (4.1). For the sake of concreteness, we provide an example.
Example 4.6.
Consider seven points $x_{1},\ldots,x_{7}$ forming a two-dimensional configuration $X$. So $X$ has codimension $c=4$. Let us determine a basis for $\mathbb{W}_{0}(X)$ using volumes and proposition 4.5.
Suppose $X_{4567}=\{x_{1},x_{2},x_{3}\}$ is a core of $X$; the corresponding volume $\Delta_{4567}$ is, thus, nonzero. $X_{4567}$ is contained in four Dziobek subconfigurations: $X_{567}$, $X_{467}$, $X_{457}$, $X_{456}$. These Dziobek subconfigurations produce the respective weight vectors
$W_{4}$ $=(-\Delta_{1567},\Delta_{2567},-\Delta_{3567},\Delta_{4567},0,0,0),$
$W_{5}$ $=(-\Delta_{1467},0,\Delta_{2467},-\Delta_{3467},0,\Delta_{4567},0,0),$
$W_{6}$ $=(-\Delta_{1457},\Delta_{2457},0,-\Delta_{3457},0,0,\Delta_{4567},0),$
$W_{7}$ $=(-\Delta_{1456},\Delta_{2456},-\Delta_{3456},0,0,0,\Delta_{4567}),$
where the index of each vector is the index of the point added to $X_{4567}$ in order to form the corresponding Dziobek subconfiguration. Since $\Delta_{4567}\neq 0$, clearly $W_{4},W_{5},W_{6},W_{7}$ form a basis of $\mathbb{W}_{0}(X)$.
Now, if $X_{4567}$ is not a core of $X$, then it necessarily has dimension one, and we can replace one of its points with one of the points in $X\setminus X_{4567}$, say replace $x_{3}$ with $x_{4}$. If $X_{3567}$ is a core, we construct a basis using our procedure, otherwise, we make another replacement. Such a sequence of replacements will eventually produce a core of $X$, or else $X$ is collinear, which is a contradiction. Notice that the same argument shows that replacements can be used to successively obtain each element $S_{k}$ of a Dziobek tree for $X$.
We end this section with a criterion for membership in $\mathbb{W}_{0}(X)$. The proof is followed by related remarks.
Theorem 4.7.
Let $X$ be a configuration of dimension $d$ such that $\Delta_{d+2\ldots n}\neq 0$. We have that $(w_{1},\ldots,w_{n})\in\mathbb{W}_{0}(X)$ if and only if the identities
(4.3) $(-1)^{i-d}w_{i}\Delta_{d+2\ldots n}=\sum_{l=d+2}^{n}w_{l}\Delta_{i\,d+2\ldots\widehat{l}\ldots n},\quad i=1,\ldots,d+1,$
hold, where $\widehat{l}$ indicates the removal of the index $l$.
Proof.
By hypothesis, the subconfiguration $\{x_{1},\ldots,x_{d+1}\}$ is a core of $X$. Thus we have the following basis of $\mathbb{W}_{0}(X)$:
$(-\Delta_{1d+3\,d+4\ldots n},\ldots,(-1)^{d+1}\Delta_{d+1\,d+3\,d+4\ldots n},(-1)^{d+2}\Delta_{d+2\,d+3\,d+4\ldots n},0,\ldots,0),$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \$ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ $\vdots$
$(-\Delta_{1d+2\,d+3\ldots n-1},\ldots,(-1)^{d+1}\Delta_{d+1\,d+2\,d+3\ldots n-1},0,\ldots,0,(-1)^{d+2}\Delta_{d+2\,d+3\,d+4\ldots n}),$
Each of the above weight vectors corresponds to a Dziobek subconfiguration $\{x_{1},\ldots,x_{d+1},$ $x_{l}\}$, with $l=d+2,\ldots,n$ ; we denote the weight vector associated to $l$ by $W_{l}$.
If $W=(w_{1},\ldots,\ldots,w_{n})\in\mathbb{W}_{0}(X),$ we have that $W=\sum_{d+2}^{n}a_{l}W_{l}$, with
$a_{l}=\frac{w_{l}}{(-1)^{d+2}\Delta_{d+2\ldots n}},\quad l=d+2,\ldots,n.$
On the other hand, for $i=1,\ldots,d+1$, we have that
(4.4) $w_{i}=\sum_{l=d+2}^{n}a_{l}(-1)^{i}\Delta_{i\,d+2\ldots\widehat{l}\ldots n}=\sum_{l=d+2}^{n}(-1)^{i-d}w_{l}\frac{\Delta_{i\,d+2\ldots\widehat{l}\ldots n}}{\Delta_{d+2\ldots n}},$
and the second equality promptly implies equations (4.3).
For the converse, given a weight vector $(w_{1},\ldots,w_{n})$, if we choose $a_{l}$ so that
$w_{l}=(-1)^{d+2}a_{l}\Delta_{d+2\ldots n},\ \ \ l=d+2,\ldots,n,$
then equations (4.3) imply that
$\Delta_{d+2\ldots n}\left(w_{i}-\sum_{l=d+2}^{n}(-1)^{i}a_{l}\Delta_{i\,d+2\ldots\widehat{l}\ldots n}\right)=0,\ \ \ i=1,\ldots,d+1.$
Since $\Delta_{d+2\ldots n}\neq 0$, we obtain
$w_{i}=\sum_{l=d+2}^{n}(-1)^{i}a_{l}\Delta_{i\,d+2\ldots\widehat{l}\ldots n},\ \ \ i=1,\ldots,d+1.$
Hence $(w_{1},\ldots,w_{n})=\sum_{d+2}^{n}a_{l}W_{l}$, that is, $W\in\mathbb{W}_{0}(X).$
Remark 4.8.
The first implication in theorem 4.7 does not require the hypothesis $\Delta_{d+2\ldots n}\neq 0$. We illustrate the argument, which can be found in [14, 28], in the context of central configurations.
If $(w_{1},\ldots,w_{n})\in\mathbb{W}_{0}(X)$, then
$\sum_{j\neq k}w_{j}\overrightarrow{x_{k}x_{j}}=\overrightarrow{O},\quad k=1,\ldots,n.$
Set $k=1$ and take the exterior product with $\overrightarrow{x_{1}x_{3}}\wedge\ldots\wedge\overrightarrow{x_{1}x_{d+1}}$. Let $\overrightarrow{e}_{1},\ldots,\overrightarrow{e}_{d}$ be the canonical basis of $\mathbb{R}^{d}$. From
$\Delta_{d+2\ldots n}\,\overrightarrow{e}_{1}\wedge\ldots\overrightarrow{e}_{d}$ $=\overrightarrow{x_{1}x_{2}}\wedge\overrightarrow{x_{1}x_{3}}\wedge\ldots\wedge\overrightarrow{x_{1}x_{d+1}},\ \ \text{and}$
$\Delta_{2\,d+2\ldots\widehat{l}\ldots n}\,\overrightarrow{e}_{1}\wedge\ldots\overrightarrow{e}_{d}$ $=(-1)^{1-d}\overrightarrow{x_{1}x_{l}}\wedge\overrightarrow{x_{1}x_{3}}\wedge\ldots\wedge\overrightarrow{x_{1}x_{d+1}},\ \ \text{if }\ l>d+1,$
we get
$(-1)^{2-d}w_{2}\Delta_{d+2\ldots n}=\sum_{l=d+2}^{n}w_{l}\Delta_{2\,d+2\ldots\widehat{l}\ldots n},$
which is (4.3) for $i=2$.
Remark 4.9.
The choice of the core used in theorem 4.7 was just a matter of convenience. We could have chosen any other core of $X$ and determined formulæ  corresponding to (4.3) by following the same procedure. We also notice that equations (4.3) can be obtained using theorem 4.2, albeit not in as straightforward a way.
Remark 4.10.
Consider a Dziobek subconfiguration which does not contain the core used in theorem 4.7. Interestingly, if we write equations (4.3) for the weight vector associated to this Dziobek subconfiguration, we obtain identities corresponding to the Plücker relations used in Algebraic Geometry to describe Grassmannians as projective varieties [24, 25]. For instance, if $n=5$ and $d=2$, consider the two-dimensional subspace spanned by the rows of the $2\times 4$ matrix whose columns are $\overrightarrow{x_{1}x_{2}},\ldots,\overrightarrow{x_{1}x_{5}}$. The Plücker coordinates of this subspace are $p_{12}=\Delta_{23}$, $p_{13}=\Delta_{35}$, $p_{14}=\Delta_{34}$, $p_{23}=\Delta_{25}$, $p_{24}=\Delta_{24}$ and $p_{34}=\Delta_{45}$. If $X$ has core $\{x_{1},x_{2},x_{3}\}$, then $\Delta_{45}\neq 0$, and we can consider the basis of $\mathbb{W}_{0}(X)$ formed by
$W_{4}=(-\Delta_{14},\Delta_{24},-\Delta_{34},0,\Delta_{45}),\quad W_{5}=(-\Delta_{15},\Delta_{25},-\Delta_{35},\Delta_{45},0).$
If we take the weight vector associated to the Dziobek subconfiguration $\{x_{1},x_{2},x_{4},$ $x_{5}\}$, which is $(-\Delta_{13},\Delta_{23},0,-\Delta_{34},$ $\Delta_{35})$, and apply formula (4.3) with $i=2$, we find that
$\Delta_{23}\Delta_{45}-\Delta_{24}\Delta_{35}+\Delta_{25}\Delta_{34}=0,$
which corresponds to the well known unique Plücker relation $p_{12}p_{34}-p_{13}p_{24}+p_{14}p_{23}=0$ defining the Grassmannian of two-dimensional subspaces of $\mathbb{R}^{4}$ as a variety in $\mathbb{R}P^{5}$.
@@ 5. Some Applications to central configurations
We call a configuration of particles (point masses, bodies) in $\mathbb{R}^{N}$ any weighted system $(X,m)$ where $m$ is a nonvanishing function which we traditionally refer to as mass. Recall the notation used in sections 3 and 4. The points of $X$ are denoted by $x_{i}$, $m_{i}$ is the value of $m$ at $x_{i}$, $i=1,\ldots,n$, the dimension of $X$ is $d$ and the codimension of $X$ is c=(n1)dc=(n-1)-d. We may interpret $m$ as the weight vector $(m_{1},\ldots,m_{n})\in\mathbb{W}=\mathbb{R}^{n}$. Assuming the particles interact under some homogeneous potential, the corresponding acceleration vectors have the form
$\overrightarrow{\gamma}_{j}=-\sum_{i\neq j}m_{i}s_{ij}^{a}\overrightarrow{x_{i}x_{j}},\quad s_{ij}=\overrightarrow{x_{i}x_{j}}^{2},\ \ a\in\mathbb{R},$
@@
Albouy has shown that, if $a=0$, then any configuration is central, and if $a\neq 0$ and d=n1d=n-1, i.e, if $c=0$, then $X$ is a regular simplex for all values of $\mu_{0}$. In the cases $\mu_{0}\neq 0$, this assertion follows from lemma 4.1 and theorem 5.2 below (as in [4], we tacitly assume that $\lambda/\mu_{0}>0$).
@@
Multiplying both sides of (5.1) by $m_{j}$ and summing over $j$, we obtain an equation relating the constants $\lambda,x_{O},\overrightarrow{\gamma}_{O}$ and the total mass $\mu_{0}$, namely
@@
    $X$ has codimension $c\geq 1$ and $\mu_{2}$ is constant. In particular, we have the Leibniz identity $\sum_{i<j}m_{i}m_{j}s_{ij}=0$, and
@@
    If $c=1$ then, for every $j=1,\ldots,n$, $X_{j}$ is a $d$-dimensional simplex (i.e., $X_{j}$ is a core of $X$) and $x_{j}$ is not on the boundary of the convex hull of $X_{j}$ (so $X$ is a generic configuration).
  • (4)
    $X$ has dimension $d\leq n-2$, and if $\Delta_{d+2\ldots n}\neq 0$, then
    $(-1)^{i-d}m_{i}\Delta_{d+2\ldots n}=\sum_{l=d+2}^{n}m_{l}\Delta_{i\,d+2\ldots\widehat{l}\ldots n},\quad i=1,\ldots,d+1,$
@@
Part (1) follows from equation (5.2). (2) is a direct consequence of proposition 2.3, because $m$ is nonvanishing. In order to prove (3), note that $\mu_{1}=\overrightarrow{O}$ provides a nontrivial linear relation for the elements of $X$. Hence $c\geq 1$. Moreover, $\mu_{2}$ is constant according to corollary 2.6. The equations follow from corollaries 2.5 and 2.7. The assertion on the case $c=1$ follows directly from (1), (2), and the barycentric interpretation of the (nonzero) coordinates of the mass vector. Compare with proposition 4.5. Finally, (4) is immediate from part (1) and theorem 4.7. ∎
In the case $n=4$, a full classification of a special family of central configurations with $\mu_{0}=0$, including examples of central configurations with $\lambda\neq 0$, can be found in [11].
Next, let us consider the case $\mu_{0}\neq 0$. As proved in [4], $(X,m)$ is a central configuration if and only if $\overrightarrow{\gamma}_{O}$ is the null vector and $x_{O}=\text{bar}(X,m)$ is the center of mass of $(X,m)$. The central configuration equations can thus be put in the form
@@
Therefore we can apply the theory developed in sections 2 and 4 to the study of central configurations with $\mu_{0}\neq 0$. In particular, we recover and extend some known results. Additional results on central configurations are provided in section 6 and appendix B.

### v4

@@ 5. Constraints for Mutual Distances
As in section 4, let $X=\{x_{1},\ldots,x_{n}\}$. Denote the dimension of the affine closure of $X$ by $\text{dim}(X)$, or just $d$, and call it the dimension of $X$. The difference $c$ between the maximum possible value for $\text{dim}(X)$, which is n1n-1, and $d$, that is c=(n1)dc=(n-1)-d, is called the codimension of $X$, and we write $\text{codim}(X)=c$.
In the classical paper [8], Cayley established determinantal conditions for a configuration to have codimension one, and for it being co-spherical. Cayley’s conditions will be easily deduced using the matrix framework from section 4. The main result of this section is a theorem which exhibits a complete set of independent constraints on mutual distances to ensure that a configuration has a given dimension. Constraints are also discussed in appendix A, where we show how to obtain simpler, non-determinantal constraints using the HLK theorem.
@@
The augmented configuration matrix $\mathcal{X}_{a}$ has the same rank as the matrix
@@
that is, $d+1$. The nullity of $\mathcal{X}_{a}$ is thus n(d+1)=(n1)d=cn-(d+1)=(n-1)-d=c. This is the dimension of $\mathbb{W}_{0}(X)$. ∎
Notice that every configuration of codimension $c=0$ is a simplex, and so is co-spherical. Corollary 4.2 implies the following necessary condition for co-sphericity when the codimension is positive.
Corollary 5.2.
If $X$ has codimension $c>0$ and is co-spherical, then $\dim\text{\emph{K}er}\,\,\mathcal{B}(X)\geq c$ and, in particular, $\det\mathcal{B}(X)=0$.
Cayley deduced the condition $\det\mathcal{B}(X)=0$ in [8]. He showed that, in the case of four points on a circle, that is, if $n=4$ and $c=1$, this condition implies the classical Ptolemy identity. It turns out that $\det\mathcal{B}(X)=0$ is also a sufficient condition for co-sphericity (cf. [6], chapter 9).
Next we study the dimension of $X$. Borrowing the terminology from Celestial Mechanics, a configuration of codimension one will be designated as Dziobek configuration, or Dziobek subconfiguration, when it is a proper subset of a larger configuration. The statements in the following paragraph appear in similar form in reference [11].
Theorem 4.1 and lemma 5.1 imply that $\mathcal{C}(X)$ has nullity equal to $c$. Thus, for $c=0$, the rank of $\mathcal{C}(X)$ is $n+1$, so $\det\mathcal{C}(X)\neq 0$. This means that there are no constraints on mutual distances, as expected. For configurations of codimension $c>0$, $\mathcal{C}(X)$ has rank nc+1n-c+1, so all subdeterminants of $\det\mathcal{C}(X)$ of size $(n-c+2)\times(n-c+2)$ must vanish. In particular, the $(n-c+2)\times(n-c+2)$ principal minors of $\mathcal{C}(X)$, which are precisely the Cayley-Menger determinants of the Dziobek subconfigurations of $X$, are all equal to zero. We shall soon check that not all Cayley-Menger determinants of Dziobek subconfigurations are independent and that, in general, a much smaller number of such determinants suffices to guarantee the correct codimension for $X$.
We firstly provide a heuristic argument for the number of constraints that must be satisfied by the mutual distances of a configuration of dimension $d$ formed by $n$ points in Euclidean space. Let $E(d)$ be the group of Euclidean (rigid) motions in $\mathbb{E}^{d}$. The manifold of congruence classes of configurations of $n$ points in $\mathbb{E}^{d}$ modulo isometries is $\mathbb{E}^{dn}\big/E(d)$, which has dimension dnd(d+1)/2dn-d(d+1)/2. The natural coordinates to describe the congruence classes are the $n\choose 2$ mutual distances between the points. For configurations of codimension c=(n1)d>0c=(n-1)-d>0, the number of mutual distances exceeds the dimension of the manifold of congruence classes by
(5.1) ${n\choose 2}-\left[dn-\frac{d(d+1)}{2}\right]={c+1\choose 2}$,
Now we supply a formal statement and proof.
@@
${c+1\choose 2}+n-d={c+2\choose 2}$
constraints, as claimed. ∎
It should be noted that, in the proof of theorem 5.3, it is possible to choose the subconfiguration of $X\cup\{p\}$ formed by the $d$ points of $X$ and $p$ as $d$-dimensional simplexes, so the additional contraints are expressed as Cayley-Menger determinants of Dziobek subconfigurations of $X\cup\{p\}$.
@@ 6. Some Applications to central configurations
We call a configuration of point masses, or bodies, in $\mathbb{E}^{N}$, any weighted system $(X,m)$ such that $m$ is a nonvanishing function traditionally referred to as mass. As in the previous section, we denote the points of $X$ by $x_{i}$ and $m_{i}$ is the value of $m$ at $x_{i}$, $i=1,\ldots,n$. We may interpret $m$ as the weight vector $(m_{1},\ldots,m_{n})\in\mathbb{W}(X)=\mathbb{R}^{n}$. Assuming the bodies interact under a homogeneous potential, the corresponding acceleration vectors have the form
$\overrightarrow{\gamma}_{j}=\sum_{i\neq j}m_{i}s_{ij}^{a}\overrightarrow{x_{i}x_{j}},\quad s_{ij}=\overrightarrow{x_{i}x_{j}}^{2},\ \ a\in\mathbb{R},$
@@
Albouy has shown that, if $a=0$, then any configuration is central, and if $a\neq 0$ and d=n1d=n-1, i.e, if $c=0$, then $X$ is a regular simplex for all values of the total mass $\mu_{0}$. In the cases $\mu_{0}\neq 0$, this assertion follows directly from lemma 5.1 and theorem 6.2 below (as in [3], we tacitly assume that $\lambda/\mu_{0}>0$).
@@
Multiplying both sides of (6.1) by $m_{j}$ and summing over $j$, we obtain an equation relating the constants $\lambda,x_{O},\overrightarrow{\gamma}_{O}$ and the total mass, namely
@@
    $X$ has codimension $c\geq 1$ and $\mu_{2}$ is constant. In particular, we have the Leibniz identity $\sum_{i<j}m_{i}m_{j}s_{ij}=0$, and the equations
@@
    If $c=1$ then, for every $j=1,\ldots,n$, $X_{j}$ is a $d$-dimensional simplex and $x_{j}$ is not on the boundary of the convex hull of $X_{j}$.
@@
Part (1) follows from equation (6.2), and (2) is a direct consequence of corollary 2.5, because $m$ is nonvanishing. In order to prove (3), note that $\mu_{1}=\overrightarrow{O}$ provides a nontrivial linear relation for the elements of $X$. Hence $c\geq 1$. Moreover, $\mu_{2}$ is constant according to corollary 2.7. The equations follow from corollaries 2.6 and 2.8. The assertion on the case $c=1$ follows directly from (2) and the barycentric interpretation of the (nonzero) coordinates of the mass vector. ∎
For $n=4$, a full classification of a special family of central configurations with $\mu_{0}=0$, including examples of central configurations with $\lambda\neq 0$, can be found in [9].
Next, let us suppose that that the total mass is nonzero. As proved in [3], $(X,m)$ is a central configuration if and only if $\overrightarrow{\gamma}_{O}$ is the null vector and $x_{O}=\text{bar}(X,m)$ is the center of mass of $(X,m)$. The central configuration equations can thus be put in the form
@@
Therefore we can apply the results of sections 3 and 5 to the study of central configurations with $\mu_{0}\neq 0$. In particular, we recover the important result below.